# Random Forests as Statistical Procedures: Design, Variance, and Dependence

### v2

@@ 7 Discussion
This paper develops a finite-sample, design-based characterization of random forest predictors by treating each tree as a randomized prediction rule acting on a fixed covariate configuration. Conditional on $X$, randomness arises jointly from the outcome realization $(Y\mid X)$ and the tree-generating mechanism, and prediction variability is governed by the induced distribution of terminal regions and associated weights. The resulting framework establishes random forests as explicit statistical procedures whose behavior can be analyzed through the same conditioning logic used in classical fixed-design inference.
A key motivation for this perspective is that random forests differ fundamentally from deterministic prediction procedures. For linear or generalized linear models, the prediction at a fixed point $x$ is an algebraic function of the observed data, and variability arises exclusively from sampling. In random forests, the prediction depends on both the data and the procedural randomization $\theta$, introducing a design-induced variability component that has no analogue in deterministic methods. Conditional on $X$, this additional component decomposes into Monte Carlo noise, which diminishes under sufficient aggregation, and a structural covariance floor $C_{T}(x)$, which does not. The covariance floor is a cost of randomization, paid in exchange for the benefits of ensemble diversity and the ability to trade resolution against stability through hyper-parameters.
The exact variance identity in Theorem 1 makes this separation precise. The Monte Carlo component $\sigma_{T}^{2}(x)/B$ governs seed-to-seed variability and goes to zero as $B$ grows, while $C_{T}(x)$ remains as the irreducible variance floor under infinite aggregation conditional on $X$. Decomposing $C_{T}(x)$ via the law of total covariance reveals two structurally distinct mechanisms: observation reuse, whereby shared training outcomes induce correlated predictions, and partition alignment, whereby independently generated trees discover similar local geometry near $x$ and define similar averaging rules even when their training data are disjoint. These mechanisms operate through different features of the tree-generating process and respond to different design parameters – observation reuse to the training fraction $p_{\mathrm{obs}}$, and alignment to the candidate-set size $q$ and the level of split-level randomization.
These results carry direct practical implications. Stabilization of forest predictions across random seeds for tree generation – a standard assessment of sufficient aggregation – reflects only the elimination of the Monte Carlo component and does not indicate the absence or magnitude of $C_{T}(x)$. A forest whose predictions no longer vary across constructions has converged to its infinite-aggregation target $f_{\infty}(x)=\mathbb{E}[T_{\theta}(x)\mid X]$, but the precision of that target is governed by the structural covariance between trees. This covariance floor is not detectable through repeated forest construction at fixed $B$; it is a property of the dependence structure among tree-level predictions and can only be assessed by examining their joint behavior within a single forest.
More broadly, existing asymptotic confidence intervals for random forest predictions, including those based on U-statistic and influence-function representations (Mentch and Hooker, 2016; Wager and Athey, 2018), target the sampling variability of the infinite-aggregation functional $f_{\infty}(x)$ under repeated data generation. That is, they quantify how $f_{\infty}(x)$ would vary across new datasets drawn from the same population. Within this sampling-based estimand, these procedures are correctly calibrated. However, they do not incorporate the additional design-induced variability captured by $C_{T}(x)$, which reflects dependence among trees within a fixed dataset and persists under infinite aggregation at fixed sample size. The sampling component decreases at rate $O(n^{-1})$ under standard stability conditions, while $C_{T}(x)$ at fixed design parameters need not decay at the same rate. When $C_{T}(x)$ is non-negligible relative to the sampling variance, intervals based solely on asymptotic sampling theory quantify sampling uncertainty but may not reflect the total predictive variability of the finite-sample forest procedure.
Within this framework, the effect of added randomization admits a precise finite-sample explanation. Additional randomization disperses the distribution of induced partitions, weakening alignment across trees and reducing the corresponding contribution to $C_{T}(x)$, while simultaneously altering the resolution of the induced predictor. Finite-sample improvements arise when the reduction in alignment-induced dependence outweighs the accompanying change in resolution. This mechanism follows directly from the variance identity and the covariance decomposition, without appeal to asymptotic arguments or heuristic notions of regularization. The framework thereby unifies several existing explanations for the empirical benefits of randomization in tree ensembles (Mentch and Zhou, 2020, 2022; Liu and Mazumder, 2025; Curth et al., 2024) by identifying alignment-induced dependence and resolution as the two competing design-driven mechanisms through which added randomization acts.
The present framework complements the existing theoretical literature on random forests. Prior work has established consistency, convergence rates, and asymptotic distributions under increasing sample size, typically by representing forests as weighted nearest-neighbor or U-statistic–type estimators. These results characterize sampling variability but do not isolate the finite-sample dependence induced by the tree-generating mechanism at fixed data. By conditioning on the observed covariate configuration and analyzing the design-induced distribution, the present approach identifies the structural sources of dependence that govern finite-sample behavior and persist under aggregation. Within this formulation, the object of study is the random forest predictor itself, in the same sense that one studies the properties of a fitted generalized linear model independently of how cross-validation is used to assess its performance. Out-of-bag predictions and related resampling-based performance estimates correspond to distinct estimands and should not be conflated with the predictor. Understanding how such estimators inherit and interact with the structural dependence identified here is a natural direction for future work. The covariance floor $C_{T}(x)$ is itself directly estimable from the sample covariance of tree-level predictions within a fitted forest, since these predictions are exchangeable under the design distribution conditional on $X$. Decomposing $C_{T}(x)$ into its overlap and alignment components, however, requires the conditional variance quantities $\sigma_{i}^{2}=\operatorname{Var}(Y_{i}\mid X)$, which are properties of the outcome model rather than the randomized design. Developing estimators for these individual components, together with their finite-sample properties, is deferred to subsequent work.
@@ 8 Conclusion
This paper develops a finite-sample, design-based characterization of random forest predictors by viewing each tree as a randomized tree-level prediction rule acting on a fixed covariate configuration. Conditioning on $X$ yields an exact variance identity that separates aggregation variability from structural dependence induced by the tree-generating mechanism and decomposes this dependence into observation-overlap and partition-alignment components. This framework clarifies the role of randomization in controlling finite-sample prediction variability by weakening structural dependence and reshaping resolution, and provides a principled basis for understanding how forest design parameters govern ensemble behavior beyond aggregation effects alone.
Appendix A Sampling variability of the infinite-forest target
@@
Let $\mathcal{D}_{n}^{\star}=\{Z_{i}=(X_{i},Y_{i})\}_{i=1}^{n}$ denote an i.i.d. sample from an underlying population. Assume that $f_{\infty}(x;\mathcal{D}_{n}^{\star})$ admits an asymptotically linear expansion of the form
@@
where $\mathbb{E}[\psi_{x}(Z_{i})]=0$, $\mathbb{E}[\psi_{x}(Z_{i})^{2}]<\infty$, and $\sqrt{n}\,r_{n}(x)\to 0$ in $L^{2}$.
@@
The remainder terms satisfy $\operatorname{Var}(r_{n}(x))=o(n^{-1})$ and $\operatorname{Cov}(n^{-1}\sum_{i}\psi_{x}(Z_{i}),r_{n}(x))=o(n^{-1})$ by the assumed $L^{2}$ rate.
Consequently,
@@
Such asymptotic linear representations and variance calculations are standard for stable estimators; see Mentch and Hooker (2016) and Wager and Athey (2018) for applications to random forest–type procedures.
Appendix B Proof of Lemma 1
Proof.
$S^{(m)}_{\theta,v}=\{\pi_{\theta,v}(1),\dots,\pi_{\theta,v}(m)\}.$
Under this coupling, for any $m_{1}<m_{2}$,
[…]
follows a hypergeometric distribution with expectation
Appendix C Proof of Theorem 2
@@
Appendix D Proof of Proposition 2
@@
Appendix E Proof of Proposition 3
Assumption (monotone tree growth in $s$). Trees grown under different minimum-terminal-occupancy constraints can be constructed on a common probability space such that, for fixed underlying randomization (observation selection, candidate restriction, and tie-breaking), the partition generated under $s^{\prime}<s$ is a refinement of the partition generated under $s$. Along every root-to-leaf path, the $s^{\prime}$-tree performs all splits performed by the $s$-tree and possibly additional splits, and it never reverses an earlier split.
@@ Appendix E Proof of Proposition 3
Fix a prediction point $x$ and construct two trees $\theta^{(s)}$ and $\theta^{(s^{\prime})}$ on the same probability space satisfying the monotone-growth assumption. By refinement,
$\mathcal{R}_{\theta^{(s^{\prime})}}(x)\subseteq\mathcal{R}_{\theta^{(s)}}(x)\quad\text{almost surely}.$
Because terminal-region size is defined by the number of observations contained in the region,
$\bigl|\mathcal{R}_{\theta^{(s^{\prime})}}(x)\bigr|\leq\bigl|\mathcal{R}_{\theta^{(s)}}(x)\bigr|\quad\text{almost surely}.$
Taking expectations yields

### v3

@@ 9 Discussion
This paper reframes random forests as randomized statistical procedures acting on a fixed set of covariates, and explicitly targets the variance of $\hat{f}(x)\mid X$, akin to that of standard regression procedures. This structural reframing reveals that the ensemble’s finite-sample variance is governed by two structurally distinct dependence mechanisms: observation reuse and partition alignment. Alignment is the more fundamental—it arises whenever trees trained on data from the same data-generating process discover similar local prediction rules, and persists even when sample splitting eliminates observation overlap entirely. Intuitively, this occurs because predictions are based on outcomes drawn from the same conditional parent population, governed by similar splits in the constructed trees. The covariance floor $C_{T}(x)$ that these mechanisms produce persists regardless of the number of trees in a forest and represents an intrinsic property of the deployed procedure.
The PASR estimator provides a practical framework for estimating this variance. It is unbiased for the covariance floor under the fitted model at any finite aggregation level, requires no asymptotic machinery, and degrades gracefully under difficult nuisance estimation conditions. For continuous outcomes, the cross-fitted residual product construction guarantees that any systematic error is conservative; for binary outcomes, the absence of a separate variance estimation step yields asymptotic unbiasedness at rate $O(n^{-2})$ (Proposition 4), confirmed empirically across all design configurations tested. The operational variance decomposition combines the PASR floor estimate with the within-forest Monte Carlo variance to produce prediction and confidence intervals that achieve near-nominal coverage for both outcome types, providing — for the the first time — pointwise uncertainty quantification for predicted probabilities from deployed classification forests.
The covariance floor addresses a different inferential question than existing asymptotic approaches. Methods based on U-statistic or influence-function representations (Mentch and Hooker, 2016; Wager and Athey, 2018) target the sampling variability of $f_{\infty}(x)$ under repeated data generation. The PASR-based intervals instead condition on the realized training data and quantify the total uncertainty of the deployed prediction, including design-induced variability that persists under infinite aggregation. These two perspectives are complementary rather than competing; a complete characterization of predictive uncertainty would require both.
The framework applies in high-dimensional settings without modification to its theoretical properties, since the estimator operates on i.i.d. synthetic replicates conditional on $X$. The binding constraint is nuisance estimation quality: for continuous outcomes, reliable estimation of the conditional variance function degrades when $p$ is large relative to $n$, while for binary outcomes the $O(n^{-2})$ bias rate of Proposition 4 ensures that the estimator remains accurate even at $p>n$.. Developing nuisance strategies that remain stable in high dimensions is a natural direction for future work and represents the primary obstacle to high-dimensional application, rather than any limitation of the PASR framework itself. Nevertheless, the current implementation provides usable results in high-dimensional settings—conservative but well-calibrated—particularly for the prediction of probabilities from classification forests.
Extensions to other forest procedures. In Appendix D we describe how the framework extends to quantile regression forests, random survival forests, honest forests, and causal forests. Full development and empirical validation of PASR within these procedures is deferred to future work; this discussion serves to demonstrate that the theory developed here is not specific to the standard random forest but applies broadly to tree-based ensembles whose trees are generated by an exchangeable mechanism. In each case, the tree-level prediction is a deterministic functional of realized outcomes in a random averaging set, and the variance identity in Theorem 1 applies without modification. The covariance floor is strictly positive whenever the alignment mechanism is active, regardless of the outcome type, loss function, or within-node estimator. Honest forests provide a particularly informative case: the sample-splitting design eliminates observation overlap while leaving alignment intact, confirming that alignment is the more fundamental source of dependence. The framework does not extend directly to boosted ensembles such as gradient boosting, where trees are constructed sequentially with each tree depending on the residuals of its predecessors, violating the exchangeability assumption that underlies Theorem 1. Whether analogous variance decompositions can be developed under sequential dependence is an open question for future work.
Future directions. The framework developed here opens several avenues for further investigation. The error decomposition in (7) provides a precise target for improving finite-sample performance: advances in calibration strategies or alternative nuisance estimation would tighten the nuisance gap and improve the PASR estimates for continuous outcomes. Deriving the asymptotic distribution of $\widehat{C}_{T}(x)$—a central limit theorem for the PASR estimator—would enable formal inference on the floor itself. The synthetic replicates generated by the PASR procedure contain distributional information beyond the first two moments; leveraging empirical quantiles from these replicates for coverage calibration, and exploring connections with conformal prediction methods, are natural extensions of the current interval construction. The extensions in Appendix D establish that the framework applies to causal, honest, survival, and quantile regression forests; developing tailored PASR specifications and empirical validation for these procedures would broaden the practical reach of the methodology.
Honest forests provide a particularly valuable setting for extending this framework beyond pointwise intervals to hypothesis testing and simultaneous inference. By eliminating observation overlap through sample splitting, honesty isolates the alignment mechanism as the sole source of $C_{T}(x)$, making design-conditional inference procedures more tractable. Natural extensions include testing treatment effect heterogeneity ($H_{0}:\tau(x_{1})=\tau(x_{2})$) using $\text{Cov}(\hat{\tau}(x_{1}),\hat{\tau}(x_{2})\mid X)$, subgroup identification with explicit uncertainty bounds that distinguish confident responders from uncertain regions, and simultaneous inference calibrated to the joint distribution available from PASR replicates. These applications would leverage the design-conditional variance decomposition to enable inference tasks not currently addressable with existing asymptotic sampling-based methods.
Finally, the design-based perspective suggests that forest hyperparameters could be tuned with explicit awareness of the resolution–dependence trade-off characterized in Section 6, rather than through prediction error alone.
@@ 10 Conclusion
This paper develops a finite-sample, design-based theory for random forest predictors. An exact variance identity separates aggregation noise from structural dependence, and a covariance decomposition identifies observation reuse and partition alignment as the two mechanisms generating the covariance floor. The floor is estimable via procedure-aligned synthetic resampling, and the resulting estimator is unbiased under the fitted model at any finite aggregation level. Operational prediction and confidence intervals constructed from the PASR variance decomposition achieve nominal coverage for both continuous and binary outcomes. For continuous outcomes, the nuisance gap is provably conservative; for classification forests, the PASR estimator is asymptotically unbiased at rate $O(n^{-2})$, providing—for the first time—theoretically grounded pointwise uncertainty quantification for predicted probabilities from deployed classification forests. Together, these results offer a principled basis for understanding how forest design parameters govern ensemble behavior, for quantifying the total predictive uncertainty of fitted forests, and for extending these ideas to tree-based ensembles with exchangeable tree-generating mechanisms.
Appendix A Proofs and Derivations
A.1 Sampling variability of the infinite-forest target
@@
Let $\mathcal{D}_{n}^{\star}=\{Z_{i}=(X_{i},Y_{i})\}_{i=1}^{n}$ denote an i.i.d. sample from an underlying population. Mentch and Hooker (2016) and Wager and Athey (2018) establish that, under regularity conditions, $f_{\infty}(x;\mathcal{D}_{n}^{\star})$ admits an asymptotically linear expansion of the form
@@
where $\psi_{x}$ is a mean-zero influence function with finite variance and $r_{n}(x)$ is a remainder whose variance is of smaller order than $n^{-1}$.
@@
The remainder terms satisfy $\operatorname{Var}(r_{n}(x))=o(n^{-1})$, and the cross-term is $o(n^{-1})$ by the Cauchy–Schwarz inequality applied to the covariance. Consequently,
@@
A.2 Proof of Theorem 2
@@
A.3 Proof of Lemma 1
Proof.
$S^{(m)}_{\theta,v}=\{\pi_{\theta,v}(1),\dots,\pi_{\theta,v}(m)\}.$
Under this coupling, for any $m_{1}<m_{2}$,
[…]
follows a hypergeometric distribution with expectation
A.4 Proof of Proposition 1
@@
A.5 Proof of Lemma 2
@@
Fix a prediction point $x$ and condition throughout on the realized covariate configuration $X$. By definition,
$C_{T}(x)=\operatorname{Cov}\!\bigl(T_{\theta}(x),T_{\theta^{\prime}}(x)\mid X\bigr),$
where $\theta$ and $\theta^{\prime}$ are independent draws from the tree-generating mechanism. Apply the law of total covariance with respect to $Y$: